Suppose a firm has an average total cost function given by: ATC = 2916/q +9 + 9q.

Calculate the price where the firm breaks even
Answer #1

Firm break evens when P=minimum of ATC.

To calculate the level of q at which ATC is at its minimum , take the derivative of ATC w.r.t q and equate it to 0 , we get:

-2916/q2 + 9 = 0

q2 = 2916/9

q2 = 324

q= 18

So, At break even : P=2916/18 + 9 + 9(18)

= 162 +9+ 162

P= 333 (Price where the firm break evens)